### v3

# Vibration-based damage detection of a trainer jet via multiple input tangential interpolation

Abstract
Control engineering is a highly developed field, which includes similarly advanced areas like system identification. In structural dynamics, system identification methods are employed for the extraction of modal parameters, such as natural frequencies and mode shapes, from any structure. In turn, these are the main building blocks of vibration-based damage detection. However, traditional comparisons of these parameters are often ambiguous in complex systems, complicating damage detection and assessment. The modified total modal assurance criterion (MTMAC), an index well-known in the field of finite element model updating, is extended to address this challenge and is proposed as a index for damage identification and severity assessment. To support the requirement for precise and robust modal identification of Structural Health Monitoring (SHM), the improved Loewner Framework (iLF), known for its reliability and computational performance, is pioneeringly employed within SHM. Since the MTMAC is proposed solely as a damage identification and severity assessment index, the coordinate modal assurance criterion (COMAC), also a well-established tool, but for damage localisation using mode shapes, is used for completeness. The iLF SHM capabilities are validated through comparisons with traditional methods, including least-squares complex exponential (LSCE) and stochastic subspace identification with canonical variate analysis (SSI-CVA) on a numerical case study of a cantilever beam. Furthermore, the MTMAC is validated against the traditional vibration-based approach, which involves directly comparing natural frequencies and mode shapes. Finally, an experimental dataset from a BAE Systems Hawk T1A trainer jet ground vibration tests is used to demonstrate the iLF and MTMAC capabilities on a real-life, real-size SHM problem, showing their effectiveness in detecting and assessing damage.
keywords:
Loewner Framework , System Identification , Multi-input multi-output , Experimental Modal Analysis , Structural Health Monitoring , Aeronautical Systems , MTMAC , Tangential Interpolation
[…]
[label1]organization=Department of Aerospace Engineering, Universidad Carlos III de Madrid, addressline=Av.da de la Universidad 30, postcode=28911, city=Leganés, state=Madrid, country=Spain
[…]
[label2]organization=Department of Structural, Building and Geotechnical Engineering, Politecnico di Torino,addressline=Corso Duca degli Abruzzi 24, city=Turin, postcode=10129, state=Piedmont, country=Italy
[…]
[label3]organization=Electronic Technology Department, Universidad Carlos III de Madrid, addressline=Av.da de la Universidad 30, postcode=28911, city=Leganés, state=Madrid, country=Spain
{highlights}
A system identification-driven structural health monitoring (SHM) method is proposed;
The two-step method involves modal parameter extraction and evaluation via a combined metric;
The improved Loewner Framework is used to extract the modal parameters, a first for SHM;
The modified total modal assurance criterion is used, as a first, as a standalone SHM index;
The method is validated on a numerical system and applied to a full-size real-life trainer jet.
1 Introduction
In the general control engineering domain, the field of system identification is critical as a way to extract information and/or obtain models from simulation, experimental, or real data [1]. For example, in structural dynamics, system identification methods are used for the extraction of modal parameters [2], such as natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$), from linear time invariant systems (LTIs) [3]. Modal parameters have always been the backbone of vibration-based Structural Health Monitoring (SHM) [4]. Hence, there is a continuous search for novel and efficient methods for extracting modal parameters, such as those presented in [5], alongside more refined indices [6] and analysis methods [7]. The former are necessary to move away from direct comparison of modal parameters -– a notoriously challenging task for higher-dimensional systems. For this reason, indices such as the modal assurance criterion (MAC) [8], which helps condense information to make it usable for SHM, have achieved considerable success over the years. On the other hand, the latter is not restricted to direct methods and can include indirect methods, such as model updating-based approaches [9, 10], as well as data-driven techniques [11]. This continuous stream of emerging methods and strategies is also fuelled by the fact that most engineering fields require some sort of damage monitoring or identification strategy. Some of the more well-known applications are infrastructures, such as roads [12] and wind turbines [13], and systems, such as fuel pumps, [14].
Damage can be defined as any structural change in a system that compromises or affects its operational capacity [15]]. The process of implementing a damage detection strategy for aerospace [16], civil [11], or mechanical engineering [17] systems is commonly referred to as SHM. In the case of vibration-based SHM, modal parameters, such as $\omega_{n}$, $\zeta_{n}$, and $\bm{\phi}_{n}$, are frequently selected as damage indicators due to their ease of identification from vibrational data [18] and their direct relationship with the mass and stiffness of the structure. These parameters can serve various purposes, such as assessing damage severity and/or localisation. $\omega_{n}$ are particularly effective for evaluating damage severity, but less reliable for localisation since small frequency changes are often masked by environmental and operational variations [19]. Nevertheless, they were successfully applied in [20] for the SHM of a masonry bridge structure using single input and multi output (SIMO) data. Conversely, $\bm{\phi}_{n}$ are better suited for damage localisation, typically achieved by detecting deviations between baseline and damaged values. Still, they can also be helpful for damage severity assessment. For example, [21] proposed a two-step procedure for damage detection on beam-like structures using $\bm{\phi}_{n}$. $\zeta_{n}$ are hardly ever used as a standalone damage indicator due to their strong dependence on non-structural factors, which makes it difficult to distinguish changes in damping caused by damage from those due to other influences [22]. Recent advances in vibration-based SHM have focused on civil infrastructure, particularly bridges [23] and wind turbines [24], with a growing interest in addressing challenges such as uncertainty quantification, noise robustness, and operational variability [25]. Advanced methodologies, such as the Cepstrum-Informed Attention-Based Network, have been introduced to model structural dynamics efficiently while maintaining physical interpretability, using compact damage-sensitive features derived from cepstral coefficients [26]. Bayesian modal identification approaches have been developed to incorporate ambient and seismic response data, enabling robust modal parameter estimation with reduced computational complexity [27]. Additionally, digital twin frameworks have established themselves as effective tools for predictive maintenance, offering unambiguous correlations between dynamic behaviours and structural conditions, such as under-foundation scour in bridges [28]. For wind turbine rotors, operational modal analysis techniques have been adapted to identify localised stiffness reductions and anisotropic effects using time-periodic approximations and Gaussian residual tests [29]. Furthermore, machine learning methods, including convolutional neural networks, have been successfully applied to classify damage scenarios in lightweight bridge structures, addressing the complexities of coupled dynamic responses from moving loads [30]. Lastly, population-based SHM has demonstrated the potential to transfer structural behaviour knowledge from homogeneous building populations to individual targets, offering a scalable solution for damage detection across nominally identical structures [31].
Nevertheless, an unequivocal solution for SHM does not yet exist, as most methods and strategies are either underdeveloped or not fully codified in national or international standards [32]. In addition, developing new methods – or refining existing ones – is a challenge due to the lack of diverse datasets from the system experimental setup perspective. In fact, most case studies and benchmarks for vibration-based SHM focus on single-input multi-output (SIMO) [33] or output-only data [34] rather than multiple input and multiple output (MIMO). This is critical as MIMO testing can offer further insight into modes not fully identified from SIMO or OMA campaigns, such as out-of-plane dominant motion in flexible wings [3] or civil infrastructure. A well-known example is the Z24 bridge case study, where the weakly excited lateral modes are the most difficult to identify [35]. Despite this difficulty, [36] showed that the first lateral mode was the most indicative of damage occurrence, which can be extrapolated as an argument to use identification of small amplitude vibration modes to enhance the precision of vibration-based SHM. While some efforts were made at the beginning of the 2000s for MIMO-based bridge monitoring [37, 38], these were not followed by further advances or testing campaigns. Additionally, there is a general lack of aeronautically relevant examples for SHM, with the notable exception of [39, 40]. However, this issue was recently addressed by the Laboratory for Verification and Validation (LVV) at the University of Sheffield, which conducted a thorough vibration testing campaign for a full-scale trainer jet aircraft [41] and has released the dataset [42]. This includes vibration data from different input signals, including simulated and real damaged cases in a MIMO configuration, offering a thorough standard for future identification and damage detection methods. Addressing the fact that damage assessment and localisation through modal parameters direct comparison for complex structures is difficult at best. Thus, a two-step modal & damage identification framework is proposed. The first step uses the improved Loewner Framework (iLF) [43] to identify the modal parameters in the system, while the second step uses the modified total modal assurance criterion (MTMAC) [44] as a damage assessment index based on the identified modal parameters. The iLF, a recently developed computationally efficient MIMO modal parameter identification technique, is applied in this article to the trainer jet aircraft dataset mentioned above and, beforehand, validated on a numerical case study. The work reported here represents the first instance in the literature where iLF is applied for modal parameter-based SHM purposes. To highlight the capabilities of the iLF, the results are compared to those from other industrially relevant methods, such as least-squares complex exponential (LSCE) [45] and stochastic subspace identification with canonical variate analysis (SSI-CVA) [2]. The former is a frequency domain method (as is the iLF), while the latter is an output-only technique in the time domain. In addition, the MTMAC (modified total modal assurance criterion) is shown to be a valid index for damage assessment. It is highlighted that this criterion was used before, together with MACFLEX (modal flexibility assurance criterion) [46], but is here proposed as a standalone damage assessment index. Further, the well-known Coordinate Modal Assurance Criterion (COMAC) [47] is used for damage localisation to ensure the completeness of the damage assessment procedure. The rationale is that the COMAC identifies which degrees-of-freedom (DoFs) measurements contribute to a low value of MAC [48]; hence, a low value of COMAC corresponds to a local $\bm{\phi}_{n}$ trajectory deviation, which is a typical sign of damage. This is supported, among many other examples, by the relevant literature on damage localisation on composite laminates [49], reinforced concrete beams [50], and retrofitted unreinforced masonry walls [51].
Thus, the main contributions of this work are the following:
  1. i.
    A two-step modal & damage identification framework is proposed:
    1. a.
      (First step) The iLF is shown to be valid for SHM purposes;
    2. b.
      (Second step) The MTMAC is shown to serve as a damage assessment quantification index.
  2. ii.
    The framework (iLF and MTMAC) is applied to a comprehensive dataset obtained from a full-scale trainer jet aircraft.
The remaining sections of the paper are organised as follows:
  1. 2.
    Methodology: The Loewner Framework (LF) theory, the backbone of the iLF;
  2. 3.
    Numerical case study: A MIMO numerical case study of a cantilever beam with three-step changes in its section profile is analysed. Three damage (element stiffness reduction) and a mass addition (on a single element) cases are used to validate the use of the iLF and the MTMAC;
  3. 4.
    Experimental case study: Modal parameters from twelve cases (one healthy and 11 damage scenarios) of the BAE Hawk T1A vibration dataset are extracted via the iLF and used for damage assessment via the MTMAC and damage localisation via the COMAC;
  4. 5.
    Conclusions: Final remarks and a summary of the results are presented.
2 Methodology
The following subsections outline the iLF theoretical background (Section 2.1) and the SHM tools considered (Section 2.2). Then, in Section 2.3, the proposed approach of using the iLF as a valid tool for SHM purposed is discussed
2.1 The Loewner Framework
The theoretical origin of the LF can be traced back to model order reduction methods [52, 53]. While its initial application was for modelling multi-port circuits in Electrical Engineering [54, 55], since then, it has been extended to investigateunsteady aerodynamics in Aeronautical Engineering [56] and recently applied to the extraction of modal parameters from SIMO mechanical systems for SHM [3]. The LF computational efficiency for SI via modal parameters extraction has been recently examined in [57]. The basic approach was then extended to the ground vibration testing of wings [58]. More recent work has focused on expanding the LF to output-only systems [59] and to MIMO systems, while also increasing its computational efficiency [43]; thus, introducing the improved LF, or iLF, which is the variant considered in this work. In essence, the LF approach fits a suitable transfer function to a set of complex-valued data, such as Frequency Response Functions (FRFs), by splitting the data into two subsets (or directions), from which rational tangential interpolants are built by exploiting the Loewner matrix $\mathbb{L}$, first introduced by Karl Loewner in the 1930s [60].
Before moving forward, a brief technical presentation of the LF, on which the iLF computational implementation is based, is given, starting with the description of the Loewner matrix $\mathbb{L}$. For a row array of pairs of complex numbers ($\mu_{j}$,${v}_{j}$), $j=1$,…,$q$, and a column array of pairs of complex numbers ($\lambda_{i}$,${w}_{j}$), $i=1$,…,$\rho$, such that $\lambda_{i}$, $\mu_{j}$ distinct, the associated $\bm{\mathbb{L}}$ can be defined:
$\bm{\mathbb{L}}=\begin{bmatrix}\frac{\mathbf{v}_{1}-\mathbf{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\mathbf{v}_{1}-\mathbf{w}_{\rho}}{\mu_{1}-\lambda_{\rho}}\\ \vdots&\ddots&\vdots\\ \frac{\mathbf{v}_{q}-\mathbf{w}_{1}}{\mu_{q}-\lambda_{1}}&\cdots&\frac{\mathbf{v}_{q}-\mathbf{w}_{\rho}}{\mu_{q}-\lambda_{\rho}}\\ \end{bmatrix}\>\in\mathbb{C}^{q\times\rho}$ (1)
This is valid if a known underlying function $\boldsymbol{\phi}$ exists, such that $\mathbf{w}_{i}=\boldsymbol{\phi}(\lambda_{i})$ and $\mathbf{v}_{j}=\boldsymbol{\phi}(\mu_{j}).$
Thus, considering a LTI system $\mathbf{\Sigma}$ with $m$ inputs and $n$ outputs, and $k$ internal variables, described by:
$\begin{split}\mathbf{\Sigma}:\;\mathbf{E}\frac{d}{dt}\mathbf{x}(t)=\mathbf{A}\mathbf{x}(t)+\mathbf{B}\mathbf{u}(t)\\ \mathbf{y}(t)=\mathbf{C}\mathbf{x}(t)+\mathbf{D}\mathbf{u}(t)\end{split}$ (2)
where $\mathbf{x}(t)\>\in\>\mathbb{R}^{k}$ is the state, $\mathbf{u}(t)\>\in\>\mathbb{R}^{m}$ is the input and $\mathbf{y}(t)\>\in\>\mathbb{R}^{p}$ is the output. The system matrices are $\mathbf{E}$, $\mathbf{A}\in\mathbb{R}^{k\times k}$, $\mathbf{B}\in\mathbb{R}^{k\times m}$, $\mathbf{C}\in\mathbb{R}^{p\times k}$, and $\mathbf{D}\in\mathbb{R}^{p\times m}$. Its transfer function $\mathbf{H}(s)$ can be defined as:
$\mathbf{H}(s)=\mathbf{C}(s\mathbf{E}-\mathbf{A})^{-1}\mathbf{B}+\mathbf{D}$ (3)
Next, the structure of tangential interpolation, often referred to as the rational interpolation in tangential directions, can be exploited to yield the following associated right (Equation 4) and left (Equation 5) interpolation data:
$\begin{gathered}(\lambda_{i};\mathbf{r}_{i},\mathbf{w}_{i}),\>i=1,\dots,\rho\quad\quad\begin{matrix}\mathbf{\Lambda}=\text{diag}[\lambda_{1},\dotsc,\lambda_{k}]\in\mathbb{C}^{\rho\times\rho}\\ \mathbf{R}=[\mathbf{r}_{1}\;\dotsc\mathbf{r}_{k}]\in\mathbb{C}^{m\times\rho}\\ \mathbf{W}=[\mathbf{w}_{1}\;\dotsc\;\mathbf{w}_{k}]\in\mathbb{C}^{\rho\times\rho}\end{matrix}\Bigg{\}}\end{gathered}$ (4)
Similarly, the left interpolation data:
$\begin{gathered}(\mu_{j},\mathbf{l}_{j},\mathbf{v}_{j}),\>j=1,\dots,q\quad\quad\begin{matrix}\mathbf{M}=\mathrm{diag}[\mu_{1},\dotsc,\mu_{q}]\in\mathbb{C}^{q\times q}\\ \mathbf{L}^{T}=[\mathbf{l}_{1}\;\dotsc\mathbf{l}_{q}]\in\mathbb{C}^{p\times q}\\ \mathbf{V}^{T}=[\mathbf{v}_{1}\;\dotsc\;\mathbf{v}_{q}]\in\mathbb{C}^{m\times q}\end{matrix}\Bigg{\}}\end{gathered}$ (5)
The indices $\lambda_{i}$ and $\mu_{j}$ represent, respectively, the values at which the right ($\mathbf{R}$) and left ($\mathbf{L}$) subsets are sampled to produce the approximation $\mathbf{H}(s)$, while $\rho$ and $q$ correspond to their dimension. The vectors and denote the right and left tangential general directions, typically chosen randomly in practice [57], while the values and represent the right and left tangential data.
The rational interpolation problem is resolved when the transfer function $\mathbf{H}(s)$, associated with realisation $\mathbf{\Sigma}$ in Equation 2, is linked to $\mathbf{w}_{i}$ and $\mathbf{v}_{i}$ in such a way that the Loewner pencil satisfies Equation 6:
$\begin{split}\mathbf{H}(\lambda_{i})\mathbf{r}_{i}=\mathbf{w}_{i},\>j=1.\dots,\rho\\ \mathbf{l}_{i}\mathbf{H}(\mu_{j})=\mathbf{v}_{j},\>i=1,\dots,q\end{split}$ (6)
such that the Loewner pencil satisfies Equation 6.
Now, given a set of points $Z=\{z_{1},\dots,z_{N}\}$ in the complex plane, the corresponding value of $\mathbf{H}(s)$ can be separated into left $\lambda_{i=1...\rho}$ and right $\mu_{i=1...q}$ data with $N=q+\rho$:
$\begin{split}Z=\{\lambda_{1},\dots,\lambda_{\rho}\}\cup\{\mu_{1},\dots,\mu_{q}\}\\ Y=\{\mathbf{w}_{1},\dots,\mathbf{w}_{\rho}\}\cup\{\mathbf{v}_{1},\dots,\mathbf{v}_{q}\}\end{split}$ (7)
Hence, the matrix $\bm{\mathbb{L}}$ becomes:
$\bm{\mathbb{L}}=\begin{bmatrix}\frac{\mathbf{v}_{1}\mathbf{r}_{1}-\mathbf{l}_{1}\mathbf{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\mathbf{v}_{1}\mathbf{r}\rho-\mathbf{l}_{1}\mathbf{w}\rho}{\mu_{1}-\lambda\rho}\\ \vdots&\ddots&\vdots\\ \frac{\mathbf{v}_{q}\mathbf{r}_{1}-\mathbf{l}_{q}\mathbf{w}_{1}}{\mu_{q}-\lambda_{1}}&\cdots&\frac{\mathbf{v}_{q}\mathbf{r}\rho-\mathbf{l}_{q}\mathbf{w}\rho}{\mu_{q}-\lambda\rho}\\ \end{bmatrix}\>\in\mathbb{C}^{q\times\rho}$ (8)
Since $\mathbf{v}_{q}\mathbf{r}_{\rho}$ and $\mathbf{l}_{q}\mathbf{w}_{\rho}$ are scalars, the Sylvester equation is satisfied by $\bm{\mathbb{L}}$ in such a fashion:
$\bm{\mathbb{L}}\mathbf{\Lambda}-\mathbf{M}\bm{\mathbb{L}}=\mathbf{L}\mathbf{W}-\mathbf{V}\mathbf{R}$ (9)
Thus, the shifted Loewner matrix, $\bm{\mathbb{L}}_{s}$, that is the $\bm{\mathbb{L}}$ corresponding to $s\mathbf{H}(s)$, can be defined as:
$\bm{\mathbb{L}}_{s}=\begin{bmatrix}\frac{\mu_{1}\mathbf{v}_{1}\mathbf{r}_{1}-\lambda_{1}\mathbf{l}_{1}\mathbf{w}_{1}}{\mu_{1}-\lambda_{1}}&\cdots&\frac{\mu_{1}\mathbf{v}_{1}\mathbf{r}_{\rho}-\lambda_{\rho}\mathbf{l}_{1}\mathbf{w}_{\rho}}{\mu_{1}-\lambda_{\rho}}\\ \vdots&\ddots&\vdots\\ \frac{\mu_{q}\mathbf{v}_{v}\mathbf{r}_{1}-\lambda_{1}\mathbf{l}_{q}\mathbf{w}_{1}}{\mu_{q}-\lambda_{1}}&\cdots&\frac{\mathbf{v}_{q}\mathbf{r}_{\rho}-\mathbf{l}_{q}\mathbf{w}_{\rho}}{\mu_{q}-\lambda_{\rho}}\\ \end{bmatrix}\>\in\mathbb{C}^{q\times\rho}$ (10)
Similarly, the Sylvester equation can be fulfilled now as follows:
$\bm{\mathbb{L}}_{s}\Lambda-\mathbf{M}\bm{\mathbb{L}}_{s}=\mathbf{L}\mathbf{W}\mathbf{\Lambda}-\mathbf{M}\mathbf{V}\mathbf{R}$ (11)
Without loss of generality, $\mathbf{D}$ can be considered 0 since its contribution does not affect the tangential interpolation of LF [52]. Thus, for ease of presentation, the remainder of the presentation will focus on:
$\mathbf{H}(s)=\mathbf{C}(s\mathbf{E}-\mathbf{A})^{-1}\mathbf{B}$ (12)
A realisation of minimal dimension is achievable only when the system is fully controllable and observable. Thus, if the data is assumed to be sampled from a system where the transfer function is represented by Equation 12, the generalised tangential observability, $\mathcal{O}_{q}$, and generalised tangential controllability matrices $\mathcal{R}_{\rho}$ can be derived from Equations 11 and 12 [55]: as follows:
$\mathcal{O}_{q}=\begin{bmatrix}\mathbf{l}_{1}\mathbf{C}(\mu_{1}\mathbf{E}-\mathbf{A})^{-1}\\ \vdots\\ \mathbf{l}_{q}\mathbf{C}(\mu_{q}\mathbf{E}-\mathbf{A})^{-1}\end{bmatrix}\>\in\mathbb{R}^{q\times k}$ (13)
$\mathcal{R}_{\rho}=\begin{bmatrix}(\lambda_{1}\mathbf{E}-\mathbf{A})^{-1}\mathbf{Br}_{1}&\cdots&(\lambda_{\rho}\mathbf{E}-\mathbf{A})^{-1}\mathbf{Br}_{\rho}\end{bmatrix}\>\in\mathbb{R}^{k\times\rho}$ (14)
Now, combining Equations 13 and 14 with, respectively, Equations 8 and 10 yields:
$\begin{split}\bm{\mathbb{L}}_{j,i}=\dfrac{\mathbf{v}_{j}\mathbf{r}_{i}-\mathbf{l}_{j}\mathbf{w}_{i}}{\mu_{j}-\lambda_{i}}=\dfrac{\mathbf{l}_{j}\mathbf{H}(\mu_{i})\mathbf{r}_{i}-\mathbf{l}_{j}\mathbf{H}(\lambda_{i})\mathbf{r}_{i}}{\mu_{j}-\lambda_{i}}=\\ =-\mathbf{l}_{j}\mathbf{C}(\mu_{j}\mathbf{E-\mathbf{A}})^{-1}\mathbf{E}(\lambda_{i}\mathbf{E}-\mathbf{A})^{-1}\mathbf{Br}_{i}\end{split}$ (15)
$\begin{split}(\bm{\mathbb{L}}_{s})_{j,i}=\dfrac{\mu_{j}\mathbf{v}_{j}-\lambda_{i}\mathbf{l}\mathbf{w}_{i}}{\mu_{j}-\lambda_{i}}=\dfrac{\mu_{j}\mathbf{l}_{j}\mathbf{H}(\mu_{i})\mathbf{r}_{i}-\lambda_{i}\mathbf{l}_{j}\mathbf{H}(\lambda_{i})\mathbf{r}_{i}}{\mu_{j}-\lambda_{i}}=\\ =-\mathbf{l}_{j}\mathbf{C}(\mu_{j}\mathbf{E}-\mathbf{A})^{-1}\mathbf{A}(\lambda_{i}\mathbf{E}-\mathbf{A})^{-1}\mathbf{Br}_{i}\end{split}$ (16)
Thus, for minimal data, where the right and left data have equal length ($p=v$), the following is required:
$\bm{\mathbb{L}}=-\mathcal{O}_{q}\mathbf{E}\mathcal{R}_{\rho}$ $\bm{\mathbb{L}}_{s}=-\mathcal{O}_{q}\mathbf{A}\mathcal{R}_{\rho}$ (17)
This leads to the following system matrix assignment::
$\mathbf{E}=-\bm{\mathbb{L}},$ $\mathbf{A}=-\bm{\mathbb{L}}_{s},$ $\mathbf{B}=\mathbf{V},$ $\mathbf{C}=\mathbf{W}$ (18)
Consequently, the interpolating rational function can be expressed now as:
$\mathbf{H}(s)=\mathbf{W}(\bm{\mathbb{L}}_{s}-s\bm{\mathbb{L}})^{-1}\mathbf{V}$ (19)
The previous derivation applies to the minimal data case, which is hardly ever the case when working with real-world data. However, the LF approach can be extended to redundant data. To begin, assume:
$\begin{split}\mathrm{rank}[\zeta\bm{\mathbb{L}}-\bm{\mathbb{L}}_{s}]=\mathrm{rank}[\bm{\mathbb{L}}\>\bm{\mathbb{L}}_{s}]=\\ =\mathrm{rank}\begin{bmatrix}\bm{\mathbb{L}}\\ \bm{\mathbb{L}}_{s}\end{bmatrix}=k,\;\forall\;\zeta\in\{\lambda_{j}\}\cup\{\mu_{i}\}\end{split}$ (20)
Then, the short Singular Value Decomposition (SVD) of $\zeta\bm{\mathbb{L}}-\bm{\mathbb{L}}_{s}$ is computed:
$\textrm{svd}(\zeta\bm{\mathbb{L}}-\bm{\mathbb{L}}_{s})=\mathbf{Y}\mathbf{\Sigma}_{l}\mathbf{X}$ (21)
where $\mathrm{rank}(\zeta\bm{\mathbb{L}}-\bm{\mathbb{L}}_{s})=\mathrm{rank}(\mathbf{\Sigma}_{l})=\mathrm{size}(\mathbf{\Sigma}_{l})=k,\mathbf{Y}\in\mathbb{C}^{q\times k}$ and $\mathbf{X}\in\mathbb{C}^{k\times\rho}$. Further, note that:
$\begin{split}-\mathbf{A}\mathbf{X}+\mathbf{E}\mathbf{X}\mathbf{\Sigma}_{l}=\mathbf{Y}^{*}\bm{\mathbb{L}}_{s}\mathbf{X}^{*}\mathbf{X}-\mathbf{Y}^{*}\bm{\mathbb{L}}\mathbf{X}^{*}\mathbf{X}\mathbf{\Sigma}_{l}=\\ =\mathbf{Y}^{*}(\bm{\mathbb{L}}_{s}-\bm{\mathbb{L}}\mathbf{\Sigma}_{l})=\mathbf{Y}^{*}\mathbf{V}\mathbf{R}=\mathbf{BR}\end{split}$ (22)
and likewise, $-\mathbf{Y}\mathbf{A}+\mathbf{M}\mathbf{Y}\mathbf{E}=\mathbf{L}\mathbf{C}$ such that $\mathbf{X}$ and $\mathbf{Y}$ are the generalised controllability and observability matrices for the system $\mathbf{\Sigma}$ with $\mathbf{D}=0$. After verifying the right and left interpolation conditions, the Loewner realisation for redundant data is as follows:
$\begin{split}\mathbf{E}=-\mathbf{Y}^{*}\bm{\mathbb{L}}\mathbf{X},\quad\quad\mathbf{A}=-\mathbf{Y}^{*}\bm{\mathbb{L}}_{s}\mathbf{X,}\\ \mathbf{B}=\mathbf{Y}^{*}\mathbf{V},\quad\quad\mathbf{C}=\mathbf{W}\mathbf{X}\end{split}$ (23)
Finally, the system modal parameters can be extracted through an eigen-analysis of the system matrices $\mathbf{A}$ and $\mathbf{C}$ in Equation 23. This was the main contribution of [3]. The reader is encouraged to consult [52, 61] for a more comprehensive discussion of each step of the LF, [43] for insights on the iLF computational implementation, and [62] for its MATLAB implementation.
2.2 Damage sssessment, quantification, and localisation
As hinted in Section 1, damage can be assessed by identifying a difference in $\omega_{n}$ between the baseline and a successive structural state. Furthermore, quantifying this difference can pinpoint to the severity level of the damage, e.g., higher differences might be related to a more severe damage state. On the other hand, localisation is achieved by comparing the baseline and damaged $\bm{\phi}_{n}$, where deviations from the baseline shape indicate the damage location. These methodologies are well-established in the literature. Nonetheless, these deviations can also be used for assessment and quantification since a more significant deviation corresponds to a more severe damage scenario. Nevertheless, these comparisons become challenging to visualise and interpret for complex structures. Hence, in this work, the complement of the well-known MTMAC [44] is proposed for damage assessment, to pair the $\omega_{n}$ and $\bm{\phi}_{n}$ contributions to damage assessment and quantification:
$\text{MTMAC}=1-\prod^{n}_{i=1}\dfrac{\text{MAC}(\bm{\phi}_{n}^{E},\bm{\phi}_{n}^{N})}{1+\left|\frac{\omega_{i}^{N}-\omega_{i}^{E}}{\omega_{i}^{N}+\omega_{i}^{E}}\right|}$ (24)
Note that the subscript $i$ is the mode of interest, and the superscripts $E$ and $N$ are, respectively, the baseline and damaged condition. The MAC is defined as per the classical work of Allemang & Brown [63]:
$\text{MAC}(i,j)=\dfrac{\left|(\bm{\phi}_{i}^{E})^{T}(\bm{\phi}_{j}^{N})\right|^{2}}{\left((\bm{\phi}_{i}^{E})^{T}(\bm{\phi}_{i}^{E})\right)\left((\bm{\phi}_{j}^{N})^{T}(\bm{\phi}_{j}^{N})\right)}$ (25)
where subscripts $i$ and $j$ are the two modes of interest. Since the MTMAC is a function of both $\omega_{n}$ and $\bm{\phi}_{n}$, its value diminishes proportionally with the deviation between the baseline and damaged modal parameters, due to the local variations in mass or stiffness. On the other hand, damage localisation suffers from the same issues as damage quantification. Therefore, in this work, the COMAC [47] is used for damage localisation since it has been historically used for determining the local sensitivity to system changes between two sets of $\bm{\phi}_{n}$. The COMAC is defined as:
$\text{COMAC}(p)=\dfrac{\sum_{i=1}^{n}\left|(\bm{\phi}_{p,i}^{E})(\bm{\phi}_{p,i}^{N})\right|^{2}}{\sum_{i=1}^{n}(\bm{\phi}_{p,i}^{E})^{2}\sum_{i=1}^{n}(\bm{\phi}_{p,i}^{N})^{2}}$ (26)
where $p$ represents the corresponding DoF of the system.
To benchmark the use of the MTMAC as a damage assessment and severity index and the effectiveness of the COMAC for damage localisation, two traditional damage quantification and localisation methods will also be applied to the same case study: the absolute difference of $\omega_{n}$ in percentage and $\bm{\phi}_{n}$ trajectory deviations.
2.3 A novel two-step modal & damage identification method
Noting that all SHM methods depend on identified modal parameters, an accurate and efficient SHM method can be developed by merging the previous two tools, the iLF and the MTMAC.
The full workflow for this novel SHM method is:
  1. 1.
    Modal identification of a given system;
  2. 2.
    The results of the above are used to validate the MTMAC for damage assessment and quantification;
  3. 3.
    For completeness, the COMAC is used to locate the damage.
In order to benchmark the above method, the more conventional LSCE and SSI-CVA methods are used, and the results are compared to iLF. Furthermore, traditional vibration-based damage assessment, quantification, and localisation comparing $\omega_{n}$ and $\bm{\phi}_{n}$ is used to validate the MTMAC as a damage index.
3 Numerical case study
To assess the suitability of the iLF and the MTMAC for damage detection, a numerical case study of a 2D hollow rectangular cross-section beam is introduced. The beam overall geometry cross-section is shown in Figure 1. It is made of Aluminium, with a Young’s modulus (E) of 69 GPa and density ($\rho$) of 2700 kgm-3, and a length (L) of 1.8 m. The beam is discretised, in MATLAB, into four 2D (y and z displacements) Euler-Bernoulli beam elements of equal length. Sample beam element mass and stiffness matrices are available in Appendix A.
Figure 1: Numerical case study: schematic of the 2D beam with the dimensions of the cross-section. Not in scale.
[…]
Table 1: Numerical case study: Analytical and identified $\omega_{n}$ and $\zeta_{n}$ from the Baseline case
Mode # Natural Frequency [Hz] – (difference [%]) Damping Ratio [-] – (difference [%])
Analytical iLF SSI-CVA LSCE Analytical iLF SSI-CVA LSCE
1 9.23 9.23 9.23 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
2 13.23 13.23 13.23 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
3 57.92 57.92 57.92 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
4 83.01 83.01 83.01 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
5 163.25 163.25 163.25 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
6 233.96 233.96 233.96 0.02 0.02 0.02
(0.00) (0.00) (0.00) (NA) (0.00) (0.00) (0.00) (NA)
7 322.06 322.06 322.06 322.06 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
8 461.54 461.54 461.54 461.54 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
9 599.02 599.02 599.02 599.02 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
10 858.45 858.45 858.45 858.45 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
11 962.02 962.02 962.02 962.02 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
12 1378.67 1378.67 1378.67 1378.67 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
13 1525.13 1525.13 1525.13 1525.13 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
14 2185.65 2185.65 2185.65 2185.65 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
15 2502.41 2502.41 2502.41 2502.41 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
16 3586.19 3586.19 3586.19 3586.19 0.02 0.02 0.02 0.02
(0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00) (0.00)
Figure 2: Numerical case study: Stabilisation diagrams for the modes identified via iLF (Figure 2(a)), LSCE (Figure 2(a)), and SSI-CVA (Figure 2(c)) for the baseline scenario.
Figure 3: Numerical case study: Difference between the iLF-identified $\omega_{n}$ of the baseline scenario and Cases #2-5.
The maximum error of the identified parameters ( and ) reported in Table 1 is 0%. However, LSCE fails to identify lower modes in higher-dimensional systems with closely spaced modes, such as modes 1-2, as well as modes 3-6. These results align with expectations, as this issue is well-documented in the literature [43, 3]. Conversely, Figure 2 clearly shows that the stabilisation diagrams from iLF (Figure 2(a)) and SSI-CVA (Figure 2(c)) can accurately identify the stable modes, while for LSCE (Figure 2(b)) this is troublesome, resulting in only ten (out of sixteen) modes identified as stable. It should be noted, although not included in Table 1, that the MAC between the analytical $\bm{\phi}_{n}$ and those identified via iLF and SSI-CVA is equal to, or very close to, 1. On the other hand, the LSCE identified $\bm{\phi}_{n}$ MAC, when compared to the analytical values, never exceeds 0.58.
These results demonstrate that the iLF is, for MIMO experimental modal analysis testing, (1) at least as accurate as SSI-CVA and (2) significantly more accurate than LSCE. In this instance, cases with artificially added white Gaussian noise are not considered, as it has already been demonstrated in [43] that measurement noise has a very limited effect on the accuracy of the iLF identification.
3.1 Investigation of damage effects
Given the good accuracy and precision shown for the baseline case by the iLF when compared to the SSI-CVA and LSCE methods, only the iLF-derived modal parameters are used for damage assessment in the numerical case study. The iLF capability to support damage detection will be assessed by applying it to the four other cases (cases #2-5) and determining if a relationship exists between the absolute difference of the damaged and baseline $\omega_{n}$. In addition, the $\bm{\phi}_{n}$ change across the different cases will be used for damage localisation.
Figure 3 shows the aforementioned absolute difference of natural frequencies ($\text{log}|\Delta\omega_{n}|$) between the damaged and baseline cases, while Figure 4 displays the dominant axis displacement (as marked in the plot secondary y-axis) for all $\bm{\phi}_{n}$.
From Figure 3, it is clear that a relationship between damage intensity and identified frequency is confirmed, as the difference increases with increased damage. It should be noted that despite the absolute value plot, all changes are negative, as expected for stiffness reductions and mass additions. On the other hand, a more notable change is observed for the case with the added mass, which, similarly, exhibits a clear distinction from the baseline case. In addition, the iLF captured the slight differences in the identified $\omega_{n}$, as those shown for modes #15-16, which are well under 0.5% in absolute value. Hence, it can be said that the iLF can perform $\omega_{n}$-based damage detection and damage severity assessment on a numerical system.
Figure 4: Numerical case study: $\bm{\phi}_{n}$ of the modes identified by iLF. Note that the secondary y-axis labels represent the dominant displacement shown.
Figure 4 shows the $\bm{\phi}_{n}$ identified via iLF for the baseline and damaged cases. Recalling that the damage has been simulated in the second element between nodes #2 and 3, it can be seen that the deviation increases in that area with increasing damage. However, a more precise assessment can be carried out in the zoomed-in view of $\bm{\phi}_{1,2}$ in Figure 5, where it is even more apparent that both the damage intensity and position are correctly identified via the modal parameter extraction with iLF.
Figure 5: Numerical case study: Detailed view of $\bm{\phi}_{1,2}$ identified by iLF, showing the trajectory change according to the case number. Note that the secondary y-axis labels represent the dominant displacement shown.
Now that it has been assessed that the iLF can serve as a suitable tool for damage detection, the MTMAC capabilities should be verified. The MTMAC values between the damaged and added mass cases versus the baseline case are compared in Figure 6. These are computed considering all modes identified via iLF.
Figure 6: Numerical case study: MTMAC between the baseline and cases #2-5 (damaged) modal parameters identified with iLF.
It is observed that the MTMAC increases with damage, exhibiting sufficient granularity to differentiate between 5% increments (a minimal change). This demonstrates that the MTMAC can be effectively used to assess and quantify damage (at least in the ideal conditions provided by numerical cases). In addition, the larger MTMAC value for case #5 indicates that this index can distinguish between damage cases and added mass cases that alter the structure behaviour.
Further, the COMAC is employed in Figure 7 to validate its applicability to damage localisation. A similar situation to that described in Figures 4 and 5 is observed: the COMAC values decrease significantly around nodes #2 and 4, i.e., the damaged area. In particular, for case #5, the COMAC deviation precisely detects the added mass location at node #3. Note that only the results for y-direction modal displacements are presented, for the sake of brevity, since a similar result is found for the z-direction.
Figure 7: Numerical case study: COMAC values between the iLF identified $\bm{\phi}_{n}$ of the baseline case and cases #2-5. (Figure 7(a)) Across the y-axis. (Figure 7(b)) Across the y-axis – Detail for COMAC $\in$ [0.995, 1].
4 Experimental case study
The BAE Systems Hawk T1A experimental dataset has recently been made available by the Dynamics Research Group at the University of Sheffield [41]. This dataset extends a previous database [64] that focused on the starboard wing of the same aircraft. The decommissioned full trainer jet aircraft, once in service with the British Royal Air Force, is shown in Figure 8 at the LVV facilities at the University of Sheffield. The data used in this study refer to the entire Hawk airframe structure, retrieved from [42]. The same dataset was also used previously to validate the MIMO extension of the LF for linear SI [43] and in [65] for the preliminary validation of Convolutional Neural Network with Reversed Mapping approach to SHM. Here in this work, the dataset is used to validate the proposed damage assessment procedure. For the sake of brevity, only the key aspects of the experimental setup, testing procedures, and recorded data used in this research work will be detailed. A more detailed discussion can be found in [41, 64].
Figure 8: BAE Systems Hawk T1A aircraft experimental setup at the LVV at the University of Sheffield (Retrieved from [41]).
4.1 Experimental setup
The airframe was equipped with 85 PCB Piezotronics (uniaxial and triaxial) accelerometers, featureing different nominal sensitivities ranging from 10 mVg-1 to 100 mVg-1, along with other sensors, totalling 139 recording channels. The sensor layout was designed to cover specific target components: the two wings, the two horizontal stabilisers, the vertical stabiliser, key points on the fuselage, and the three landing gears. For reasons described later, only the accelerometers located on the port wing have been considered in this study. None of the other sensors -– fibre-Bragg grating strain gauges, temperature sensors, etc. -– were considered. For MIMO linear SI, five TiraTM TV 51140-MOSP modal shakers [^2] https://www.tira-gmbh.de/fileadmin/inhalte/download/schwingprueftechnik/schwingpruefanlagen/modal/EN/100N_bis_2700N/data_sheet_system_tv_51140-MOSP_eng_V05.pdf were used, attached to PCB PiezotronicsTM 208C02 force transducers (sensitivity: 11.11 mVN-1) [^3] https://www.pcb.com/products?m=208c02, applying input forces to the respective locations as indicated in Figure 9.
Figure 9: BAE Systems Hawk T1A aircraft: Shaker positions for the MIMO tests on the starboard (Figure 9(a)) and the port wing ((Figure 9(b))), the horizontal stabiliser ((Figure 9(c))), and the vertical stabiliser (Figure 9(d)). Adapted from [41].
4.1.1 Testing procedures
The dataset contains white Gaussian noise (WGN) inputs, with five amplitude levels increasing from 0.1 V to 0.5 V. Pink Gaussian noise excitations are also available for linear SI. In contrast, odd random-phase multi-sine excitations can be used to investigate the nonlinear response of the system (which is out of the scope of this work). SIMO and MIMO tests were performed, roving the input source location for SIMO tests. However, these have not been considered here either since the LF was already proved viable for SIMO vibration-based dynamic monitoring in previous works [43].
4.1.2 Recorded data and selected recordings
The acceleration time series were recorded at all output channels with a sampling frequency of $f_{s}=$ 2048 Hz. The data used here were retrieved directly as pre-processed transfer functions, defined over 8192 spectral lines. Further, the dataset contains (as of the time it was last accessed, 14th of June 2024) 216 different tests, accounting for different types of excitations, input amplitudes, number and location(s) of the input source(s), and kind of damage (simulated –- mass addition –- or actual -– panel removal).
The criteria followed in this work for the selection of the cases were:
  1. 1.
    For MIMO testing, only tests using the five available input sources were considered;
  2. 2.
    To replicate what was done in [43] and for consistency with [64], WGN inputs were selected. These were defined by an excitation bandwidth in the 5–256 Hz range;
  3. 3.
    The maximum input amplitude voltage, 0.5 V, was considered for all cases;
  4. 4.
    Six cases of added mass were considered. Specifically, these arise from adding different weights at three locations: the tip of the leading edge (PW_TLE), the root of the same (PW_RLE), and the centre of the trailing edge (PW_CTE), with the mass always added on the upper side, as shown in Figure 10. The added weights considered are M1 (254.3 g), M2 (616.8 g), and M3 (916.8 g).
  5. 5.
    The interest was focused on damage cases simulated on the port wing, which covered several types of damage. These include removed panels (at different locations) and added masses (located at different sites and with different weights to simulate growing damage severity).
  6. 6.
    All the relevant damage scenarios were separately tested.
  7. 7.
    Five damage cases were used, each one consisting of a single removed panel, from PW1 to PW5, moving from the fuselage to the wing tip (see Figure 10). Taken together, they account for different damage locations, severity, and typologies.
[…]
Thus, a total of twelve cases were considered. These are summarised in Table 2.
Table 2: BAE Systems Hawk T1A aircraft: Test scenarios considered.
Case # Scenario Description
1 HS_WN05 Healthy state with 0.5 V input
2 DS_WN10 M1 mass addition at PW_CTE
3 DS_WN12 M3 mass addition at PW_CTE
4 DS_WN13 M1 mass addition at PW_RLE
5 DS_WN15 M3 mass addition at PW_RLE
6 DS_WN17 M2 mass addition at PW_TLE
7 DS_WN18 M3 mass addition at PW_TLE
8 PR_WN03 PW1 panel removal
9 PR_WN06 PW2 panel removal
10 PR_WN09 PW3 panel removal
11 PR_WN12 PW4 panel removal
12 PR_WN15 PW5 panel removal
Moreover, aiming at MIMO SI and vibration-based SHM, these signals have all been compared to the health baseline for the same input amplitude (HS_WN5), using the 21 uniaxial 10 mVg-1 accelerometers included on the wing port. These sensors were recording in the vertical direction only, arranged in a grid of two lines on the top side (running parallel to the leading and trailing edges) and one line on the bottom side (parallel to the leading edge), with respectively of 14 and 7 uniaxial accelerometers, for a total of 21 output channels (see Figure 11).
Figure 11: BAE Systems Hawk T1A aircraft: Sensors position (X) on the port wing (adapted from [41]). Note: In the original image of the bottom view [41], the sensors were ordered from the wing tip to the root. However, the numbering in the dataset metadata [42] reflects the modified figure presented here.
4.2 Results
Before addressing the effectiveness of the proposed methodology for SHM purposes, it is worthwhile to investigate the accuracy of modal identification via iLF compared to that obtained with the benchmark methods SSI-CVA and LSCE on this experimental dataset. A further source of comparison is identified in the literature, i.e. the results reported in [43]. However, please consider that those results were obtained for the case with 0.4 V input voltage, while all data in this work refer to the 0.5 V input case. The first step involves identifying the stable modes for the healthy case, case #1, via iLF and benchmark methods. In Figure 12, the three stabilisation diagrams are presented. It is clear that the iLF (Figure 12(a)) identification succeeds in finding the most stable modes, while SSI-CVA (Figure 12(c)) and especially LSCE (Figure 12(b)) identifications return fewer and less stable poles.
Figure 12: BAE Systems Hawk T1A aircraft: Stabilisation diagrams for the modes identified via iLF (Figure 12(a)), LSCE (Figure 12(a)), and SSI-CVA (Figure 12(c)) for the baseline scenario.
From [43] and the current iLF identification, 41 modes were identified. However, only the first three are used here for the SHM task. Hence, in Table 3, the first three pairs of natural frequencies ($\omega_{n}$) and damping ratios ($\bm{\phi}_{n}$) identified in this work for the undamaged Hawk T1A Aircraft are presented alongside the aforementioned literature results in Table 3, while the iLF-identified $\omega_{n}$ and $\bm{\phi}_{n}$ for all cases are shown Figures 13(a) and 13(b).
Table 3: BAE Systems Hawk T1A aircraft: $\omega_{1-3}$ and $\zeta_{1-3}$ identified via iLF, SSI-CVA, and LSCE. 0.4 V input case refers to the literature result in [43] for a healthy case with a lower input amplitude (transposed).
Method Mode 1 Mode 2 Mode 3
Natural Frequency [Hz] – (Difference [%])
0.4 V input [43] 6.98 15.43 16.32
iLF 7.03 15.42 16.28
(0.7) (-0.1) (-0.3)
SSI-CVA
LSCE 16.25
(–) (–) (-0.4)
Damping Ratio [-] – (Difference [%])
0.4 V input [43] 0.03 0.01 0.01
iLF 0.03 0.01 0.01
(0) (0) (0)
SSI-CVA
LSCE 0.03
(–) (–) (200)
Figure 13: BAE Systems Hawk T1A aircraft: $\omega_{1-3}$ (Figure 13(a)) and $\zeta_{1-3}$ (Figure 13(b)) identified via iLF for the 12 cases under scrutiny.
As shown in Figure 13, a deviation is observed between the identified $\omega_{1-3}$ of cases #2-12 and case #1. However, linking a frequency trend (decrease or increase) to the damage (mass addition or panel removal) in this complex structure is not a straightforward task. This is a well-known issue with complex assembled systems. However, this falls in the context of damage (or more correctly, anomaly) classification, which is another task, even more difficult than damage localisation and severity assessment. Damage classification is almost impossible to assess in an unsupervised fashion [66] and, therefore, is outside of the aim of this research work. Similarly, the situation shown in Figure 13 for $\zeta_{1-3}$ does not offer any cues for inferring a clear (linear or nonlinear) relationship between the damage case and the identified damping. Nevertheless, this is particularly expected due to the damping own nature [5].
Finally, the $\bm{\phi}_{1-3}$ identified via iLF for the healthy state are considered in Figure 14. It should be noted that representing these mode shapes is particularly challenging, as the wing has been isolated from the entire aircraft system, to allow for local scaling of the modal displacements. Nevertheless, the dominant displacement motion of the port wing can be recalled from [43]:
  1. Mode #1 1st bending (downwards);
  2. Mode #2 2nd bending (downwards);
  3. Mode #3 1st coupled between bending (downwards) and torsion.
Figure 14: BAE Systems Hawk T1A aircraft: $\bm{\phi}_{1-3}$ identified via iLF for case #1.
The identified $\bm{\phi}_{1-3}$ for cases #2-12 can be compared to those of case #1 via the MAC value. These are shown in Table 4. Here, it can be seen that as the mass increases (cases #2-7), the MAC values of one or more modes go down, showing a change in the $\bm{\phi}_{1-3}$ trajectory. These changes become even more pronounced for the panel-removed cases (cases #8-12), in particular for cases #8 and 11.
Table 4: BAE Systems Hawk T1A aircraft: MAC matrices diagonal values between the damaged and removed panel cases and case #1, the healthy state.
MAC Values [-] – main diagonal terms
Case Mode 1 Mode 2 Mode 3
2 0.99 1.00 0.99
3 0.92 1.00 0.97
4 0.86 0.99 0.98
5 0.72 1.00 0.93
6 0.90 0.96 0.99
7 0.85 1.00 0.98
8 0.58 0.99 0.98
9 0.80 0.99 0.99
10 0.88 0.91 0.99
11 0.60 0.91 0.99
12 0.85 0.90 0.99
Hence, we can assess that the model parameters, specifically $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ of the damaged cases, deviate from those of case #1. However, assessing such a system by simple comparison is not a straightforward task. Hence, the combined use of MTMAC (for damage severity assessment) and COMAC (for damage localisation), already validated on the numerical system, is here tested for the damage assessment of this experimental case study, as reported in the next sections.
4.2.1 Experimentally simulated damage
In this subsubsection, the simulated damage conditions (via mass addition), corresponding to cases #2-7, are investigated. In Figure 15, the MTMAC values for these cases are shown. A relationship exists between the different mass quantities and positions. In fact, cases #4 and 5 show the highest MTMAC values of the six cases. Hence, there is a more pronounced modal parameters change with the baseline case. Traditionally, this is interpreted as a location within the structure that is more sensitive to damage effects.
Figure 15: BAE Systems Hawk T1A aircraft: MTMAC for cases #2-7 w.r.t. case #1.
Now, by looking at all cases with the same mass (M3 in cases #3, 5, and 7), it can be asserted that the location PW_RLE is the most sensitive to damage and that PW_CTE is the least, while PW_TLE lays in the middle. Nevertheless, a deviation from the baseline $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ can be identified in all three locations; most importantly, the magnitude of the added mass influences them by increasing the deviations.
Given that the MTMAC can indicate the general state, in terms of damage, of the system for the given cases, the damage localisation task remains. To this aim, the scaled COMAC of cases #2-7 are plotted in Figure 16. Please note that the COMAC values are normalised between 0 and 1 to better highlight the mode shape trajectory changes, as the damage quantification is already addressed with the MTMAC. The results are first discussed for the case pairs with the same additional mass location, followed by a global discussion.
Figure 16: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases #1 (Figure 16(a)), 2 (Figure 16(b)), 3 (Figure 16(c)), 4 (Figure 16(d)), 5 (Figure 16(e)), and 6 (Figure 16(f)).
Cases #2 and 3 (Figures 16(a) and 16(b)): from the two images, it is clear that the areas most influenced by the added mass near the centre of the trailing edge (TE) are near the tip and closer to the root at $z\approx$2 m. Notably, as the mass is increased from M1 to M3, the deviation in the mode shapes becomes more dominant near the tip, while retaining some deviation at the same near root area.
Cases #4 and 5 (Figures 16(c) and 16(d)): In these cases, the most impacted areas are near the leading edge (LE) root and around mid-span. In particular, for the M1 case, the mid-span deviation is more dominant than that at the root. However, this changes for M3, as the root sees the highest deviation and the mid-span change extends, with a lower magnitude, towards the tip. As will be discussed later, this is the only case where the highest deviation exactly matches the damage location.
Cases #6 and 7 (Figures 16(e) and 16(f)): the most prominent areas where a deviation in $\bm{\phi}_{1-3}$ are seen are the near-tip region and the same close to the root section in cases #2 and 3. Notably, the maximum deviation for both cases is seen near the damaged area. However, when the mass is increased from M2 to M3, the maximum deviation point shifts slightly towards the root (from $x\approx$4 m to -3.5 m).
4.2.2 Real, panel-off, damage
In Figure 17, the MTMAC values for the panel-off cases (#8-12) are shown. Notably, the values for the panels removed at PW2-3 and 5 have a similar magnitude, while for the PW1 and PW4 cases, the MTMAC values are doubled. This larger deviations in $\omega_{1-3}$ and $\bm{\phi}_{1-3}$ from case #1 show that the latter locations have a higher sensitivity to damage.
Figure 17: BAE Systems Hawk T1A aircraft: MTMAC values for cases #8-12 w.r.t. case #1.
Figure 18 shows the scaled COMAC for all of the panel-off cases. It is important to see that for all different locations, different deviations are identified with the COMAC; however, the maximum COMAC never exactly matches the damaged area. This will be discussed later in the paper.
Figure 18: BAE Systems Hawk T1A aircraft: Scaled COMAC for the port wing identified via iLF from cases #10 (Figure 18(a)), 11 (Figure 18(b)), 12 (Figure 18(c)), 13 (Figure 18(d)), and 14 (Figure 18(e)).
5 Discussion
For all cases, added mass and panel-off, the MTMAC successfully detects a change in the system modal parameters $\omega_{1-3}$ and $\bm{\phi}_{1-3}$. In particular, from the two subsets (added mass and panel-off), it is possible to see what locations are more sensitive to changes in the structure. The highest MTMAC values are recorded for cases #8 and 11, which, respectively, represent panel-off cases at locations near the root (PW1) and at mid-span (PW4). The other three panel-off cases show a similar MTMAC value, hence showing a similar sensitivity. Notably, all panels are roughly placed between the quarter chord and mid-chord. On the other hand, the added mass locations are close to either the LE or TE. For them, the most sensitive location to damage is case #5, near the root leading edge. In contrast, the location least sensitive to damage is at TE at mid-span. Nevertheless, in all cases, it is possible to identify the increase in the added mass with an increase in the MTMAC, as shown in Figure 15.
With respect to damage localisation, this is directly achieved only for cases #5-7. Nevertheless, each couple of added mass cases is somewhat coherent in their COMAC values distribution and, most importantly, it is unequivocally different from the others. The latter also holds for the panel-off cases. These findings indicate that the COMAC can be effectively utilised for damage localisation in complex structures, particularly when paired with a high-fidelity finite element model of the system under examination or an empirical measure to correlate the COMAC value distribution with the predicted damage location. Unfortunately, neither was available for the wing under scrutiny at the time of writing; this aspect is not addressed in this work, as it lies beyond its scope, i.e., introducing the use of iLF for MTMAC-based SHM in MIMO tests.
6 Conclusions
In this work, a practical application of control, in the form of system identification for model parameter estimation, is proposed for damage detection and localisation on a complex aeronautical system. A two-step procedure, including modal parameters identification with the improved Loewner Framework (iLF) and damage detection and quantification with the modified total modal assurance criterion (MTMAC), has been introduced. This is the first instance that the iLF has been used within structural health monitoring and that the MTMAC has been employed as a standalone damage assessment index. Furthermore, as a first, damage detection and localisation have been carried out on the multi-input multi-output vibration dataset of the BAE Systems Hawk T1A aircraft from the Laboratory for Verification and Validation at the University of Sheffield. These resulted in the following findings:
  • 1.
    The iLF has been numerically and experimentally proven to detect small damage-related changes in modal parameters from multi-input multi-output (MIMO) case studies - either numerically-simulated multiple degrees of freedom systems and complex experimental data (the Hawk T1A trainer jet);
  • 2.
    The use of the MTMAC has been validated for damage assessment and quantification. In particular, it has been proved that this modal parameters-based damage-sensitive feature gives a condensed index that has been shown to work well in both numerical and experimental data;
  • 3.
    Damage was successfully identified in 11 cases of the recently introduced Hawk T1A aircraft MIMO dataset, a first in the literature.
Appendix A Euler-Bernoulli beam mass and stiffness matrices
The 2D Euler-Bernoulli beam mass and stiffness matrices are shown in, respectively, Equations 27 and 28.
$M_{e}=\frac{\rho Aa}{105}\begin{bmatrix}78&0&0&22a&27&0&0&-13a\\ 0&78&-22a&0&0&27&13a&0\\ 0&-22a&8a^{2}&0&0&-13a&-6a^{2}&0\\ 22a&0&0&8a^{2}&13a&0&0&-6a^{2}\\ 27&0&0&13a&78&0&0&-22a\\ 0&27&-13a&0&0&78&22a&0\\ 0&13a&-6a^{2}&0&0&22a&8a^{2}&0\\ -13a&0&0&-6a^{2}&-22a&0&0&8a^{2}\\ \end{bmatrix}$ (27)
$K_{e}=\begin{bmatrix}\frac{3EI_{z}}{2a^{3}}&0&0&\frac{3EI_{z}}{2a^{2}}&-\frac{3EI_{z}}{2a^{3}}&0&0&\frac{3EI_{z}}{2a^{2}}\\ 0&\frac{3EI_{y}}{2a^{3}}&-\frac{3EI_{y}}{2a^{3}}&0&0&-\frac{3EI_{y}}{2a^{2}}&-\frac{3EI_{y}}{2a^{2}}&0\\ 0&-\frac{3EI_{y}}{2a^{3}}&\frac{2EI_{y}}{a}&0&0&0&\frac{2EI_{y}}{a}&0\\ \frac{3EI_{z}}{2a^{2}}&0&0&\frac{2EI_{z}}{a}&-\frac{3EI_{z}}{2a^{2}}&0&0&\frac{EI_{z}}{a}\\ -\frac{3EI_{z}}{2a^{3}}&0&0&-\frac{3EI_{z}}{2a^{2}}&\frac{3EI_{z}}{2a^{3}}&0&0&-\frac{3EI_{z}}{2a^{2}}\\ 0&-\frac{3EI_{y}}{2a^{2}}&0&0&0&\frac{3EI_{y}}{2a^{3}}&\frac{3EI_{y}}{2a^{2}}&0\\ 0&-\frac{3EI_{y}}{2a^{2}}&\frac{2EI_{y}}{a}&0&0&\frac{3EI_{y}}{2a^{2}}&\frac{3EI_{y}}{2a^{3}}&0\\ \frac{3EI_{z}}{2a^{2}}&0&0&\frac{EI_{z}}{a}&-\frac{3EI_{z}}{2a^{2}}&0&0&\frac{2EI_{z}}{a}\end{bmatrix}$ (28)
Corresponding author
[…]
E-mail address: gdessena@ing.uc3m.es
[…]
E-mail address: marco.civera@polito.it
Author contributions
Conceptualisation, G.D., and M.C.; methodology, G.D., and M.C.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D. and M.C.; resources, G.D, M.C., A.M., B.C.; data curation, G.D. and M.C..; writing—original draft preparation, G.D. and M.C.; writing—review and editing, G.D, M.C., A.M., B.C., and OE.BM.; visualisation, G.D., M.C, and A.M.; funding acquisition, G.D, M.C., B.C., A.M., and OE.BM..
Declaration of conflicting interests
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Funding
The first author is supported by the Grants for research activity of young PhD holders, part of the Universidad Carlos III de Madrid (UC3M) Own Research Program (Ayudas para la Actividad Investigadora de los Jóvenes Doctores, del Programa Propio de Investigación de la UC3M). Furthermore, the authors from UC3M disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work has been supported by the Madrid Government (Comunidad de Madrid – Spain) under the Multiannual Agreement with the UC3M (IA_aCTRl-CM-UC3M). The second and fourth authors are supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management). The third author gladly acknowledges funding as Beatriz Galindo Distinguished Senior Researcher by the Spanish Government.
Acknowledgements
The BAE T1A Hawk Full Structure Modal Test has been retrieved from the University of Sheffield data repository ORDA and is cited in this work. The authors thank the dataset authors and the LVV at the University of Sheffield for making them openly available.
Data availability statement
The experimental and numerical results (Data file 1) supporting this work are openly available from the Zenodo repository at [https://zenodo.org/records/13866467]. In addition, this study used third-party experimental data (Data file 2) made available at [orda.shef.ac.uk/articles/dataset/BAE_T1A_Hawk_Full_Structure_Modal_Test/24948549] under licence that the authors do not have permission to share:
  • 1.
    Data file 1. Data supporting: Vibration-based damage detection of a trainer jet via multiple input tangential interpolation;
  • 2.
    Data file 2. BAE T1A Hawk Full Structure Modal Test
Data file 1 is available under the terms of the [GNU General Public License v3.0 (GPL 3.0)].